# An Overview of Prototype Formulations for Interpretable Deep Learning

### v3

@@ B.4 Fisher-Bingham Distribution
For a $D$ dimensional space, the FB distribution is by a $D\times D$ matrix ${\bm{A}}$ of orthogonal vectors $(\bm{{\alpha}}_{1},\bm{{\alpha}}_{2},\ldots,\bm{{\alpha}}_{D})$, concentration parameter $\kappa$ and ellipticity factors $[\beta]_{2:D}$ where $\sum_{j=2}^{D}\beta_{j}=1$ and $0\leq 2|\beta_{j}|<\kappa$. The FB’s unnormalized PDF is defined as
@@ B.5 Mixture Models
Appendix C Scaled Dot-Product Prototype
@@
Appendix D Implementation Details
D.1 Data Preprocessing
D.2 Hyperparameters
The prototypical part networks ProtoPNet and HyperPGNet use a convolutional neck after the feature encoder and work on a latent feature map of size $7\times 7\times 128$. The models were trained with a minibatch size of 48 images using AdamW optimizer with learning rate 1e-4 and weight decay 1e-4.
D.3 Compute Resources
All experiments were performed on a desktop workstation with a single NVIDIA RTX 3070Ti GPU (8 GB VRAM) per model. On CUB-200-2011 the training duration was roughly 40 seconds per epoch, regardless of the used prototype formulation.
D.4 Model Implementation
Appendix E Parameter Distribution of HyperPG

### v4

@@ D.4 Fisher-Bingham Distribution
For a $D$ dimensional space, the FB distribution is by a $D\times D$ matrix $\bm{A}$} of orthogonal vectors $(\bm{{\alpha}}_{1},\bm{{\alpha}}_{2},\ldots,\bm{{\alpha}}_{D})$, concentration parameter $\kappa$ and ellipticity factors $\beta]_{2:D}$ where $\sum_{j=2}^{D}\beta_{j}=1$ and $0\leq 2|\beta_{j}|<\kappa$. The FB’s unnormalized PDF is defined as
@@ D.5 Mixture Models
Appendix E Scaled Dot-Product Prototype
@@
Appendix F Parameter Distribution of HyperPG
@@
Appendix G Asset Licenses
Our work relies on the publicly available Github Code of ProtoPNet (Chen et al., 2019) which is released under the MIT license.
Our experiments use the following datasets:
  • CUB-200-2011 (Wah et al., 2011) by Caltech Vision Lab. The use is restricted to non-commercial research and educational purposes.
  • Oxford Flowers (Nilsback and Zisserman, 2008) from the Visual Geometry Group at University of Oxford. No license is detailed.
  • Stanford Cars (Krause et al., 2013) as published on Kaggle. No license is detailed.